Emily is 68 years old. Colin is 17 years older than Emily. Dan is 16 years younger than Emily. What is the total of their combined ages?

Answer by josmiceli(19441)

Emily is 68
Colin is 68 + 17 = 85
Dan is 68 - 16 = 52
------------------------
Add their ages
68 + 85 + 52 = 205